# ShiZhi: A Chinese Lightweight Large Language Model for Court View Generation

### v1

@@ Abstract
Criminal Court View Generation (CVG) is a fundamental task in legal artificial intelligence, aiming to automatically generate the "Court View" section of a legal case document. Generating court views is challenging due to the diversity and complexity of case facts, and directly generating from raw facts may limit performance. In this paper, we present ShiZhi, the first large language model (LLM) specifically designed for court view generation. We construct a Chinese Court View Generation dataset, CCVG, of more than 110K cases, each containing fact descriptions paired with corresponding court views. Based on this dataset, ShiZhi achieving 58.5 BLEU-1 on court view generation and 86.1% accuracy with 92.5% macro F1 on charge prediction. Experimental results demonstrate that even a small LLM can generate reasonable and legally coherent court views when trained on high-quality domain-specific data. Our model and dataset are available at https://github.com/ZhitianHou/ShiZhi.
@@ 3.3 Post Filtering
To ensure data quality and compatibility with our model architecture, we apply length-based filtering. We discard any sample where either the fact or court view text is shorter than 50 characters or longer than 512 characters after extraction, which helps reduce noise and computational inefficiency. The length statistic of fact and court view in train and test sets are shown in Figure 4 and their descriptive statistics presented in the Table 1.
@@
Qwen2-0.5B-Instruct 0.0 0.0 0.0 11.1 6.3 2.9 17.3 29.5
ShiZhi 6.3 0.8 6.3 58.5 51.0 41.7 86.5 92.8
@@ 4 Model Training
The model was trained for 3 epochs with a batch size of 4 and a learning rate of 1e-3. The maximum input length was set to 512 tokens. Training was conducted on a single NVIDIA RTX 4090 GPU.
@@ 5.2 Main Results
Table 2 presents the performance of ShiZhi compared to the baseline Qwen2-0.5B-Instruct on court view generation and charge prediction. For court view generation, ShiZhi achieves ROUGE-1 and ROUGE-L scores of 6.3, while the baseline scores are 0.0. In terms of BLEU, ShiZhi attains a BLEU-1 of 58.5, markedly higher than the baseline. The observed discrepancy between relatively low ROUGE scores and substantially higher BLEU scores can be attributed to the high lexical variability in court view texts. ROUGE metrics rely on exact n-gram overlap, which penalizes semantically correct outputs that use different wording from the reference. BLEU, by contrast, is more sensitive to partial n-gram matches, capturing alignment in local phrasing even when the overall wording differs.
For charge prediction, ShiZhi achieves 86.1% accuracy and 92.5% macro F1, substantially surpassing the baseline performance of 17.3% and 29.5%, respectively. These results indicate that domain-specific fine-tuning allows ShiZhi to effectively learn the mapping from fact descriptions to charges, reflecting strong legal reasoning capabilities.
@@ 6 Conclusion
In this paper, we presented ShiZhi, the first large language model specifically designed for Criminal Court View Generation (CVG). We curated CCVG, a dataset of over 110K Chinese criminal cases with fact descriptions paired with court views, and fine-tuned a 0.5B lightweight LLM on this dataset. ShiZhi achieves 58.5 BLEU-1 on court view generation and 86.1% accuracy with 92.5% macro F1 on charge prediction, demonstrating that even a small LLM can generate legally coherent and factually grounded court views when trained on high-quality domain-specific data. Our model and dataset provide a foundation for future research in automated legal document generation and legal reasoning with LLMs.

### v2

@@ Abstract
Criminal Court View Generation (CVG) is a fundamental task in legal artificial intelligence, aiming to automatically generate the "Court View" section of a legal case document. Generating court views is challenging due to the diversity and complexity of case facts, and directly generating from raw facts may limit performance. In this paper, we present ShiZhi, the first large language model (LLM) specifically designed for court view generation. We construct a Chinese Court View Generation dataset, CCVG, of more than 110K cases, each containing fact descriptions paired with corresponding court views. Based on this dataset, ShiZhi achieving 70.00 ROUGE-1 and 67.85 BLEU-1 on court view generation, as well as 86.48% accuracy with 92.75% macro F1 on charge prediction. Experimental results demonstrate that even a small LLM can generate reasonable and legally coherent court views when trained on high-quality domain-specific data. Our model and dataset are available at https://github.com/ZhitianHou/ShiZhi.
@@ 3.3 Post Filtering
To ensure data quality and compatibility with our model architecture, we apply length-based filtering. We discard any sample where either the fact or court view text is shorter than 50 characters or longer than 1024 characters after extraction, which helps reduce noise and computational inefficiency. The length statistic of fact and court view in train and test sets are shown in Figure 4 and their descriptive statistics presented in the Table 1.
@@
Qwen2-0.5B-Instruct 24.18 6.53 16.40 29.16 17.11 9.24 16.29 28.02
ShiZhi 70.00 51.07 59.61 67.85 62.92 54.76 86.48 92.75
@@ 4 Model Training
The model was trained for 3 epochs with a batch size of 4 and a learning rate of 1e-3. The maximum input length was set to 1024 tokens. Training was conducted on a single NVIDIA RTX 4090 GPU.
@@ 5.2 Main Results
Table 2 reports the performance of different models on court view generation and charge prediction. Among the compared models, ShiZhi consistently outperforms the baseline Qwen2-0.5B-Instruct across all metrics. For court view generation, ShiZhi achieves a ROUGE-1 of 70.00, ROUGE-2 of 51.07, and ROUGE-L of 59.61, substantially higher than the baseline scores of 24.18, 6.53, and 16.40, respectively. Similarly, BLEU scores are markedly improved, with BLEU-1 at 67.85 compared to 29.16 for the baseline. The large gap between ShiZhi and the baseline demonstrates the effectiveness of domain-specific training for capturing legal reasoning and structuring coherent court views.
For charge prediction, ShiZhi achieves 86.48% accuracy and 92.75% macro F1, substantially surpassing the baseline performance of 16.29% and 28.02%, respectively. These results indicate that domain-specific fine-tuning allows ShiZhi to effectively learn the mapping from fact descriptions to charges, reflecting strong legal reasoning capabilities.
@@ 6 Conclusion
In this paper, we presented ShiZhi, the first large language model specifically designed for Criminal Court View Generation (CVG). We curated CCVG, a dataset of over 110K Chinese criminal cases with fact descriptions paired with court views, and fine-tuned a 0.5B lightweight LLM on this dataset. ShiZhi achieves 70.00 ROUGE-1 and 67.85 BLEU-1 on court view generation, as well as 86.48% accuracy with 92.75% macro F1 on charge prediction, demonstrating that even a small LLM can generate legally coherent and factually grounded court views when trained on high-quality domain-specific data. Our model and dataset provide a foundation for future research in automated legal document generation and legal reasoning with LLMs.